OpenStudy (anonymous):

standard form help? y = -4/9x - 9.7

OpenStudy (raffle_snaffle):

standard form?

OpenStudy (anonymous):

here's the full question it gives me 4x = 9y = -23, and then tells me to write the equation of a line that would be parallel to that line in standard form with points (12, -15), so i came up with the slope-intercept form of a line that's parallel to that: y = mx + b y = -15 x = 12 -15 = m12 + b the slope is -4/9 so -15 = (-4/9)12 + b solve for b b = 9.7 so it'd be y = -4/9x + 9.7 right? how do i write this in standard form?

OpenStudy (raffle_snaffle):

http://www.mathsisfun.com/algebra/standard-form.html

OpenStudy (raffle_snaffle):

Set it equal to zero.

OpenStudy (anonymous):

oh wait b is wrong, -4/-9 = +.4. .4 x 12 = 4.8 4.8 + b = -15 b = -19.8?

OpenStudy (raffle_snaffle):

DId you read the link I posted above?

OpenStudy (anonymous):

oh wait b is wrong, -4/-9 = +.4. .4 x 12 = 4.8 4.8 + b = -15 b = -19.8?

OpenStudy (anonymous):

i need to make sure my slope-intercept form is right first. y = -4/9x -19.8?

OpenStudy (raffle_snaffle):

Do you know how to check it?

OpenStudy (anonymous):

-15 = -4/9(12) - 19.8 -4/9 = .4 .4 x 12 = 4.8 -15 = 4.8 - 19.8

OpenStudy (raffle_snaffle):

yep

OpenStudy (anonymous):

to write it in standard form y = -4/9x - 19.8 multiply each side by -9 -9y = -4x - (-178.2) add 4x to each side 4x - 9y = 178.2 subtract 178.2 from each side 4x - 9y - 178.2 = 0 but my book says the answer is 4x - 9y - 183 = 0. where does 183 come from? closest i can get is 178.2

OpenStudy (raffle_snaffle):

go back and make sure b is correct or your answer key could be incorrect

OpenStudy (anonymous):

b has to be correct because -4/-9 = 0.4 repeating

OpenStudy (raffle_snaffle):

Well everything looks good to me.

OpenStudy (anonymous):

thank you!